calculate distance at which the interaction energy between two opposite elementary charges is equal to thermal energy (kbt) at 300k

1. vaccum
2.in water (e=80)
1
Expert's answer
2020-10-20T08:35:47-0400
"F = k{|q_1q_2| \over r^2}"

At 300K ktb = 0.0259 eV


"k = {1 \over 4\pi \epsilon}"

"r^2 = {k|q_1q_2| \over 0.0259} = {|q_1q_2| \over 0.0259*4\pi \epsilon}"

a)


"r^2 = {|1.6*10^{-19}*(-1.6*10^{-19})| \over 0.325304*0.0885*10^{-11}}"